The formula for the perimeter of a rectangle is P=2l+2w, where l is the length and w is the width. A rectangle has a perimeter of 24 inches. Find its dimensions if its length is 3 inches greater than its width.
Answer by checkley71(8403)

L=W+3
2(W+3)+2W=24
2W+6+2W=24
4W=24-6
4W=18
W=18/4
W=4.5 ANSWER.
L=4.5+3=7.5 ANSWER.
PROOF
2*4.5+2*7.5=24
9+14=24
24=24